The sides of a rectangular lot having an area of 5.85 hectares were measured using a 100 m. tape
that was 0.025 m. too short. The length of the longer side of the lot is 3/2 of the shorter side.
____________ a.) Compute the corrected length of the longer side
____________ b.) Compute the corrected length of the shorter side
1
Expert's answer
2021-10-15T01:20:03-0400

5.85 Hectares

=0.000585m2

Let the shorter side be=b

Longer side= bX3/2

Area= LXW

0.000585=b(bX3/2)

0.000585=1.5b2

1.5b=0.024

b=0.024/1.5

b=0.02m

longer side =0.02X1.5

=0.03m